# ToMiE: Towards Modular Growth in Enhanced SMPL for 3D Human Gaussians with Animatable Garments

### v1

@@ 5.2 Baselines and Metrics
Given the wide variety of work on the human body, we select only the most cutting-edge and representative works from each focus area for a fair comparison. Since 3DGS is currently the leading representation for novel view synthesis, we compare 3DGS-based methods, including 3DGS-Avatar [37], GART [21], and GauHuman [14]. Among them, GART is expected to offer extra animatability for its modeling of implicit global auxiliary bones. Additionally, there is another category of human modeling without incorporating SMPL-like pose priors. Although these methods don’t guarantee an animatable human avatar, they can achieve high-quality rendering, among which, we compare the rendering quality of Im4D [25] with our method.
We conduct a comprehensive qualitative and quantitative comparison of our ToMiE against these methods. We report three key metrics: peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM) [41], and learned perceptual image patch similarity (LPIPS) [43]. Per-scene results can be found in the supplemental material. In addition to comparing the rendering results, we also demonstrate the animatability of the extra garments enabled by our method. We strongly recommend readers to watch the supplemental video for a more intuitive understanding of the animating.
@@ 5.3 Novel View Synthesis Results
Fig. 4 and Tab. 1 present the results of our method compared to other baselines. In the tables, we showcase two evaluation protocols. The first evaluates the entire image, reflecting the overall rendering quality. The second uses a binary mask to specifically compare the complex garments regions, demonstrating how our method outperforms others in these challenging cases. The mask is shown in Fig. 4, whose details can be found in the supplemental material.
Table 1: Quantitative comparison between our method and other runners-up. $\mathbb{B}$ and $\mathbb{G}$ stand for human body and complex garments. We color each result as best, second best and third best.
Dataset Animatable DNA-Rendering ZJU-Mocap
@@
3DGS-Avatar        27.69 0.948 0.0496 16.96 30.61 0.970 0.0296
GART        29.25 0.958 0.0480 17.12 30.91 0.962 0.0318
Im4D        26.28 0.965 0.0308 14.90 28.99 0.973 0.0620
GauHuman        30.22 0.962 0.0405 18.26 30.82 0.962 0.0326
ToMiE(ours)        31.28 0.966 0.0374 19.75 31.10 0.963 0.0312
@@ 5.4 Ablation Studies
A. Modular Growth Ablation. We remove the adaptive growth process to ablate its impact on the rendering results. Tab. 2 “w/o Modular Growth” shows a decline in rendering quality, while the human garments also become not animatable.
@@
Table 2: Ablation studies on DNA-Rendering dataset [5]. We independently ablate the modular growth strategy and non-rigid design to validate their impact on the overall performance.
@@
5.5 Animating Hand-held Objects and Loose-fitting Garments
Figure 5: Animating Results of ToMiE and GART [21]. Our explicit modeling fully decouples garments from the human body, enabling part-specific animating. However, implicit modeling of garments conditioned on human poses leads to the failure of animating in GART.
We demonstrate the uniqueness of our method, specifically its ability to explicitly animate complex garments. Our animating approach can be implemented in two ways. On the one hand, we can utilize the transformation already recorded in the extra joint book to replay garments motions. The visualization of this part is shown in Fig. 5. On the other hand, ToMiE also supports bypassing the decoding process of the extra joint book by directly inputting the transformation explicitly. This allows us to customize the motion trajectory of external joints. Due to the limitations of the image’s expressive capabilities, we strongly recommend the readers to watch the supplemental video to check the animated results under the “spiral” trajectory.
In Fig. 5, we edit the extra joints of complex garments while keeping the body poses under the SMPL paradigm stationary. Since the implicit auxiliary bones of GART [21] is controlled by the traditional SMPL poses, only the identical appearance can be output when SMPL poses are stationary. In contrast, our method explicitly models the garments, fully decoupling them from the traditional SMPL poses, enabling free animating.
@@ 6 Limitations and Conclusion
Limitations. Although our method enhances the modeling for rigid and non-rigid garments, it cannot address scenarios involving drastic changes in the topology. This is because topological changes disrupt the one-to-one correspondence between frames, making the human modeling centered on the canonical space become downgraded. We notice that [33] address topology issues by introducing high-dimensional mappings, which could be adapted to build our non-rigid deformation. However, this is not the main scope of this paper and can be explored as future work.
Conclusion. In this paper, we introduce ToMiE, a modular growth method designed to extend traditional SMPL skeleton for better modeling of human garments. In the first stage, we assign the gradient of gaussian points to different joints by combining LBS weights with the motion kernel based on motion priors. This allows us to accurately locate the parent joints that need to grow, avoiding redundant growth. In the second stage, we design an extra joint book to achieve explicit joint modeling and optimize the transformation of the newly grown joints in an end-to-end manner. With the improved designs mentioned above, our ToMiE stands out among numerous state-of-the-art methods, achieving the best rendering quality and animatability of garments. We hope the adaptive growing method will spark a renewed discussion on the current capabilities of digital human modeling. What’s more, it is expected to offer some insights for subsequent works related to topology and skeleton generation.
Appendix A Supplemental Material
A.1 Per-scene Results on DNA-Rendering Dataset
We exhibit our per-scene results on DNA-Rendering dataset [5] in Tab. 4. To more clearly demonstrate the results of the growth, we further present the parent joints $J\in\bm{J}_{s}$ for each case in Tab. 3, as defined in Sec. 4.2. We present more visualizations on the DNA-Rendering dataset in Fig. 8.
@@ A.1 Per-scene Results on DNA-Rendering Dataset
A.2 Visualization on ZJU-Mocap Dataset
In Sec. 5.1, we quantitatively experiment on ZJU-Mocap [36] dataset to validate that our method is also effective in scenarios with tight-fitting garments. In Fig. 7 , we present more visualization results.
@@ A.2 Visualization on ZJU-Mocap Dataset
Figure 8: More qualitative results on the DNA-Rendering [5] dataset. We show cases of hand-held objects (0800_07) and loose-fitting garments (0811_06) (from top to bottom).
A.3 Calculation of Masks with Complex Garments
Our method focuses primarily on modeling complex garments. In Tab. 1, we further evaluate the model’s performance in these areas using a binary mask. To generate a mask for each scene that distinguishes regions containing complex garments, we first segment the potential complex garment points in 3D space. Specifically, points predominantly controlled by the extra joints and their parent joints are identified as part of the complex garments. With the pre-trained blending weight, we can easily locate these points, which are then assigned a white color, while all others are marked in black, forming a 3D binary mask. Finally, we obtain a 2D binary mask representing complex garments by applying Gaussian rasterization to the 3D mask. Since we rely solely on this binary mask for metrics evaluation, this post-processing method for calculating masks is permissible.
A.4 Adjustment of the Gradient Threshold for Densification
In scenarios with complex garments, to prevent excessive gaussian points from causing high memory consumption, we propose an adaptive suppression strategy to keep the number of gaussian points within a reasonable range. This is achieved by dynamically adjusting the threshold of the gradient for densification $\epsilon_{d}$ in [18]. This threshold $\epsilon_{d}$ represents that points with accumulated gradients exceeding it will be densified. Therefore, a larger threshold results in fewer split points, and vice versa.
@@ A.4 Adjustment of the Gradient Threshold for Densification
A.5 Details of Hyperparameters
The $\lambda$ of balancing MK weight and LBS weight in Eq. 8 is set to $0.4$. The gradient threshold $\epsilon_{\bm{J}}$ to identify the $J\in\bm{J}_{s}$ that requires growth in Eq. 13 is set to $3.5\times 10^{-6}$. For network hyperparameters, we detail the number of layers and the width of the MLP network design. $\Phi_{\text{lbs}}$ has $D=4$ and $W=128$. $\Phi_{p}$ has $D=4$ and $W=256$. $\Phi_{r}$ has $D=4$ and $W=128$. $\Phi_{d}$ has $D=2$ and $W=128$. Specifically, we initialize the weights of the last layer in $\Phi_{p}$ and $\Phi_{r}$ as a tiny value $1\times 10^{-2}$. This provides stability for the initial training phase. Warm-up iterations number in Sec. 4.4 is set to $8\times 10^{3}$.
A.6 Supplemental Video
Our supplemental video consists of three parts. First, we present monocular rendering results to demonstrate that we can accurately render complex garments on the human body, including hand-held objects. Next, we perform 360-degree rendering of the full video to validate our generalization capability on novel views. Finally, we fix standard human skeleton still, only animating extra garments to show our decoupling and explicit animating capability. Since image quality may not fully demonstrate the effectiveness of human reconstruction, especially for animating results, we recommend that readers refer to this supplemental video for better visualization.

### v2

@@ 5.2 Baselines and Metrics
We select the most cutting-edge and representative works from each focus area for a fair comparison. Since 3DGS is currently the leading representation for novel view synthesis, we compare 3DGS-based methods, including 3DGS-Avatar [46], GART [29], and GauHuman [17]. Among them, GART is expected to offer extra animatability for its modeling of implicit global auxiliary bones. Additionally, there is another category of human modeling without incorporating SMPL-like pose priors. Although these methods don’t guarantee an animatable human avatar, they can achieve high-quality rendering, among which, we compare the rendering quality of Im4D [33] with our method.
We conduct a comprehensive comparison of our ToMiE against these methods. We report three key metrics: peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM) [51], and learned perceptual image patch similarity (LPIPS) [56]. Per-scene results can be found in the supplemental material. In addition to comparing the rendering results, we also demonstrate the animatability of the complicated human parts enabled by our method. We strongly recommend readers to watch the supplemental video for a more intuitive understanding of the animating.
@@ 5.3 Novel View Synthesis Results
Fig. 4Tab. 1 and Tab. 2 present the results of our method compared to other baselines. In the tables, we showcase two evaluation protocols. The first evaluates the entire image, reflecting the overall rendering quality. The second uses a binary mask to specifically compare the complicated regions, demonstrating how our method outperforms others in these challenging cases. The mask is shown in Fig. 4, whose details can be found in the supplemental material.
Dataset Animatable DNA-Rendering
@@
3DGS-Avatar        28.93 0.953 0.0459 17.69
GART        29.25 0.958 0.0480 17.12
Im4D        26.28 0.965 0.0308 14.90
GauHuman        30.22 0.962 0.0405 18.26
ToMiE(ours)        31.28 0.966 0.0374 19.75
Table 1: Quantitative comparison between our method and other methods on complicated DNA-Rendering dataset [7]. $\mathbb{B}$ and $\mathbb{G}$ stand for human body and complex clothing (including hand-held objects). We color each result as best, second best and third best. ToMiE achieves the best performance in PSNR and SSIM, while ranking second only to rendering-based Im4D [33] in LPIPS, mainly because rendering-based methods do not consider human structural constraints, resulting in higher visual fidelity.
Dataset Animatable ZJU-Mocap
[…]
3DGS-Avatar        30.61 0.970 0.0296
GART        30.91 0.962 0.0318
Im4D        28.99 0.973 0.0620
GauHuman        30.82 0.962 0.0326
ToMiE(ours)        31.10 0.963 0.0312
Table 2: Quantitative comparison between our method and other methods on ZJU-Mocap dataset [45] with tight-fitting clothing. $\mathbb{B}$ and $\mathbb{G}$ stand for human body and clothing (including hand-held objects). We color each result as best, second best and third best. This comparison demonstrates that ToMiE can achieve comparable (SSIM and LPIPS) or even superior (PSNR) performance to other approaches in tight-fitting scenarios where joint growth is not required.
@@ 5.4 Ablation Studies
A. Adaptive Growth Ablation. We remove the adaptive growth to ablate its impact on the rendering results. Tab. 3 “w/o Adaptive Growth” shows a decline in rendering quality, while the hand-held objects and loose-fitting clothing also become not animatable.
@@
Table 3: Ablation studies on DNA-Rendering dataset [7]. We independently ablate the adaptive growth strategy and non-rigid design to validate their impact on the overall performance.
5.5 Animating Hand-held Objects and Loose-fitting clothing
We demonstrate the uniqueness of our method, specifically its ability to explicitly animate hand-held objects and loose-fitting clothing. Our animating approach can be implemented in two ways. On the one hand, we can utilize the transformation already recorded in the extra joint book to replay clothing motions. The visualization of this part is shown in Fig. 5. On the other hand, ToMiE also supports bypassing the decoding process of the extra joint book by directly inputting the transformation explicitly. This allows us to customize the motion trajectory of external joints. Due to the limitations of the image’s expressive capabilities, we strongly recommend the readers to watch the supplemental video to check the animated results.
In Fig. 5, we edit the extra joints while keeping the body poses under the SMPL paradigm stationary. Since the implicit auxiliary bones of GART [29] is controlled by the traditional SMPL poses, only the identical appearance can be output when SMPL poses are stationary. In contrast, our method explicitly models hand-held objects and loose-fitting clothing, fully decoupling them from the traditional SMPL poses, enabling free animating.
Figure 5: Animating Results of ToMiE. Our explicit modeling fully decouples hand-held objects and loose-fitting clothing from the human body, enabling part-specific animating.
@@ 6 Limitations and Conclusion
Limitations. Although our method enhances the modeling for rigid and non-rigid clothing, it cannot address scenarios involving drastic changes in the topology (e.g., taking off clothes or opening a book). This is because topological changes disrupt the one-to-one correspondence between frames, making the human modeling centered on the canonical space become downgraded. We notice that [41] addresses topology issues by introducing high-dimensional mappings, which could be adapted to build our non-rigid deformation. However, this is not the main scope of this paper and can be explored as future work.
Conclusion. In this paper, we introduce ToMiE, an adaptive growth method designed to extend traditional SMPL skeleton for better modeling of hand-held objects and loose-fitting clothing. In the first stage, we assign the gradient of gaussian points to different joints by combining LBS weights with the motion kernel based on motion priors. This allows us to accurately locate the parent joints that need to grow, avoiding redundant growth. In the second stage, we design an extra joint book to achieve explicit joint modeling and optimize the transformation of the newly grown joints in an end-to-end manner. With the improved designs mentioned above, our ToMiE stands out among numerous state-of-the-art methods, achieving the best rendering quality and animatability of hand-held objects and loose-fitting clothing. We hope the adaptive growing method will spark a renewed discussion on the current capabilities of digital human modeling. What’s more, it is expected to offer some insights for subsequent works related to topology and skeleton generation.
Appendix A Per-scene Results on DNA-Rendering Dataset
We exhibit our per-scene results on DNA-Rendering dataset [7] in Tab. 5. To more clearly demonstrate the results of the growth, we further present the parent joints $J\in\bm{J}_{s}$ for each case in Tab. 4, as defined in Sec. 4.2 (main text). We present more visualizations on the DNA-Rendering dataset in Fig. 8.
@@ Appendix A Per-scene Results on DNA-Rendering Dataset
Appendix B Visualization on ZJU-Mocap Dataset
In Sec. 5.1 (main text), we quantitatively experiment on ZJU-Mocap [45] dataset to validate that our method is also effective in scenarios with tight-fitting garments. In Fig. 7 , we present more visualization results.
@@ Appendix B Visualization on ZJU-Mocap Dataset
Figure 8: More qualitative results on the DNA-Rendering [7] dataset. We show cases of hand-held objects (0041_10) and loose-fitting garments (0007_04) (from top to bottom).
Figure 9: Qualitative Comparison of Im4D [33].
Appendix C LBS Visualization
We visualize the LBS weights in Fig. 10, including regions corresponding to maximum and second largest skinning weights.
Figure 10: Visualization of LBS weights.
Appendix D Ablations on Motion Kernel
Our motion kernel aids in the gradient assignment of complex human bodies, helping to accurately identify the joints that need to be expanded. In Fig.3 (main text), the white mask indicates gaussian gradients that belong to the wrist (red point). With MK only, gradients between the wrist and the elbow was incorrectly counted (“misdirect”). With LBS only, the feather duster was partly lost (“misclassify”). So in Eq.(8) (main text), we integrate both to determine parent joints needing growth. We show ablations in Tab. 6 and Fig. 11.
Figure 11: Ablation on Motion Kernel and LBS.
[…]
ToMiE(MK only) 30.27 0.958 0.0432 18.33
ToMiE(LBS only) 30.41 0.960 0.0413 18.52
ToMiE(full) 31.28 0.966 0.0374 19.75
Table 6: Ablation on Motion Kernel and LBS.
Appendix E Training and Rendering Efficiency
ToMiE Gauhuman GART 3DGS-Avatar Im4D
Training Time 30min 25min 15min 20min 3h
Rendering Speed (FPS) 60+ 180+ 60+ 50+ 0.5
Table 7: Average training (till convergence) time and rendering speed on DNA-Rendering dataset [7].
Tab. 7 reports our average training time and rendering speed on DNA-Rendering dataset [7]. We train all these methods on a single GeForce RTX3090.
Appendix F Calculation of Masks with Hand-held Objects and Loose-fitting Garments
Our method focuses primarily on modeling hand-held objects and loose-fitting Garments. In Tab. 1 (main text) and Tab. 2 (main text), we further evaluate the model’s performance in these areas using a binary mask. To generate a mask for each scene that distinguishes regions containing hand-held objects and loose-fitting Garments, we first segment the potential complex garment points in 3D space. Specifically, points predominantly controlled by the extra joints and their parent joints are identified as part of the hand-held objects and loose-fitting Garments. With the pre-trained blending weight, we can easily locate these points, which are then assigned a white color, while all others are marked in black, forming a 3D binary mask. Finally, we obtain a 2D binary mask representing hand-held objects and loose-fitting Garments by applying Gaussian rasterization to the 3D mask. Since we rely solely on this binary mask for metrics evaluation, this post-processing method for calculating masks is permissible.
Appendix G Adjustment of the Gradient Threshold for Densification
In scenarios with hand-held objects and loose-fitting garments, to prevent excessive gaussian points from causing high memory consumption, we propose an adaptive suppression strategy to keep the number of gaussian points within a reasonable range. This is achieved by dynamically adjusting the threshold of the gradient for densification $\epsilon_{d}$ in [25]. This threshold $\epsilon_{d}$ represents that points with accumulated gradients exceeding it will be densified. Therefore, a larger threshold results in fewer split points, and vice versa.
@@ Appendix G Adjustment of the Gradient Threshold for Densification
Appendix H Details of Hyperparameters
The $\lambda$ of balancing MK weight and LBS weight in Eq. (8) (main text) is set to $0.4$. The gradient threshold $\epsilon_{\bm{J}}$ to identify the $J\in\bm{J}_{s}$ that requires growth in Eq. 13 is set to $3.5\times 10^{-6}$. For network hyperparameters, we detail the number of layers and the width of the MLP network design. $\Phi_{\text{lbs}}$ has $D=4$ and $W=128$. $\Phi_{p}$ has $D=4$ and $W=256$. $\Phi_{r}$ has $D=4$ and $W=128$. $\Phi_{d}$ has $D=2$ and $W=128$. Specifically, we initialize the weights of the last layer in $\Phi_{p}$ and $\Phi_{r}$ as a tiny value $1\times 10^{-2}$. This provides stability for the initial training phase. Warm-up iterations number in Sec. 4.4 (main text) is set to $8\times 10^{3}$.
Appendix I Visual Comparison with Im4D
As Im4D [33] can only perform rendering, we didn’t compare its visual quality with other animatable methods in the main text. In Fig. 9, we compare its rendering quality with our ToMiE. Although it is purely rendering-based and not constrained by SMPL skeleton, ensuring high visual fidelity (LPIPS), it suffers from the inability to animate and color distortion (PSNR).
Appendix J Supplemental Video
Our supplemental video consists of three parts. First, we present monocular rendering results to demonstrate that we can accurately render hand-held objects and loose-fitting garments on the human body. Next, we perform 360-degree rendering of the full video to validate our generalization capability on novel views. These two parts are visually compared with GauHuman [17], one of the current state-of-the-art methods. Finally, we fix standard human skeleton still, animating extra joints to show our decoupling and explicit animating capability. Since image quality may not fully demonstrate the effectiveness of human reconstruction, especially for animating results, we recommend that readers refer to this supplemental video for better visualization.